# Disturbance-Robust Backup Control Barrier Functions: Safety Under Uncertain Dynamics

### v1

@@ III Disturbance Robustness
While the (Backup CBF-QP) can guarantee safety for a dynamical system in which the dynamics are perfectly known, in practice there will always be unmodeled parameters or external disturbances which will perturb the dynamics. Therefore, it is desirable to leverage the advantages offered by the backup CBF approach, for dynamics with process disturbances. As such, consider the modified dynamical system
@@
We assume that a backup set $\mathcal{C}_{\rm B}$ and corresponding control law $\boldsymbol{u}_{\rm b}$ can be obtained which render a small subset of $\mathcal{C}_{\rm S}$ robustly forward invariant. This is made more precise below.
@@
${\dot{\boldsymbol{\phi}^{d}_{\rm b}\mkern-1.0mu}\mkern 1.0mu}{}(\tau,\boldsymbol{x})=f_{\rm cl}(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))+\boldsymbol{d}_{x},\mkern 9.0mu\boldsymbol{\phi}^{d}_{\rm b}(0,\boldsymbol{x})=\boldsymbol{x}.$ (14)
@@
Let $\mathcal{L}_{h}$ and $\mathcal{L}_{h_{\rm b}}$ be the Lipschitz constants of $h$ and $h_{\rm b}$ respectively, and let $\delta_{{\rm max}}(\tau)$ be a norm bound on the deviation between $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ and $\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})$ at time $\tau\in[0,T]$:
$\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\|\leq\delta_{{\rm max}}(\tau),$ (21)
@@
These guarantee that $h_{\rm b}(\boldsymbol{\phi}^{d}_{\rm b}(T,\boldsymbol{x}))\geq 0$. Thus, all the functions which define $\mathcal{C}_{\rm D}$ are positive, meaning that $\boldsymbol{x}\in\mathcal{C}_{\rm D}$ for all $\boldsymbol{x}\in\mathcal{C}_{\rm I}$, and so $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$. ∎
Lemma 1 assumes that a time-varying bound on the deviation between the nominal and disturbed backup flow, $\delta_{\rm max}(\tau)$, can be found. While problem-specific bounds can be obtained, we utilize a generalization of the Gronwall-Bellman inequality to obtain such a bound for a wide class of nonlinear systems.
@@
For systems (5) and (14), let $f_{\rm cl}$ be locally Lipschitz on $\mathcal{X}$ with Lipschitz constant $\mathcal{L}_{\rm cl}$ and $\boldsymbol{d}_{x}$ be piecewise continuous in $\tau$ on $[0,T]$. If $\left\|\boldsymbol{d}_{x}\right\|\leq\xi$ for all $\boldsymbol{d}_{x}$ and some $\xi>0$ then
$\left\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\right\|\leq\frac{\xi}{\mathcal{L}_{\rm cl}}(e^{\mathcal{L}_{\rm cl}\tau}-1)\triangleq\delta_{{\rm max}}(\tau),$
for all $\tau\in[0,T]$.
@@
In the case where contraction bounds on the deviation between the disturbed and nominal backup flow can be obtained, $\delta_{{\rm max}}(\tau)$ can be made much less conservative. Details on such bounds can be found in [13, Corollary 3.17]. A contraction bound is used effectively in the spacecraft rotation example in Section IV-B. For linear systems, flow deviation bounds can be made even tighter.
Using the definition of $\mathcal{C}_{\rm D}$ and the corresponding robust backup controller, we now examine the properties of $\mathcal{C}_{\rm D}$.
@@
Therefore applying definition (17) with (24) and (25) completes the proof. ∎
@@
By Lemma 1, any $\boldsymbol{x}\mkern-3.0mu\in\mkern-2.0mu\mathcal{C}_{\rm I}\mkern-6.0mu\implies\mkern-6.0mu\boldsymbol{x}\mkern-3.0mu\in\mkern-3.0mu\mathcal{C}_{\rm D}$ and by Lemma 3, $\boldsymbol{u}_{\rm b}$ is one such controller which ensures that $\boldsymbol{\phi}^{d}(\vartheta,\boldsymbol{x})\in\mathcal{C}_{\rm D}\subseteq\mathcal{C}_{\rm S},\forall\mkern 2.0mu\vartheta\geq 0$. ∎
We are now ready to apply a safety condition to obtain conditions for a controller which ensures safety of (13). From the definition of $\mathcal{C}_{\rm I}$ we have
@@
Here, $\dot{\boldsymbol{x}}^{d}\triangleq f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}+\boldsymbol{d}_{x}$. Using this expansion, we can show that a control solution which realizes forward invariance of $\mathcal{C}_{\rm I}$ keeps the disturbed system safe, and that conditions for such a controller can be directly computed, despite the unknown disturbance.
@@
From a direct application of Theorem 1 for dynamical system (13), we obtain (31). As done in [5], the robustness terms $\eta$ and $\eta_{\rm b}$ upperbound the unknown $\boldsymbol{d}_{x}$ term. The condition (36) thus ensures that for any $\boldsymbol{x}_{0}\in\mathcal{C}_{\rm I}$, $\boldsymbol{\phi}^{d}(t,\boldsymbol{x}_{0})\in\mathcal{C}_{\rm I},\forall\mkern 1.0mut>0,$ and from Lemma 1, $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$. ∎
Naturally, the original backup CBF constraints (10) are recovered in the absence of disturbances (i.e., $\xi=0$). As the constraints in (36) are continuous in $\tau$, the trajectory is again discretized and appropriately tightened via a constant term $\varepsilon_{\Delta}$ where
@@
$\begin{gathered}\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}\quad\quad\quad\quad(\text{DR-BCBF-QP})\\ \text{s.t. }\\ \nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta\geq\\ \hskip 80.00012pt\ignorespaces-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\epsilon_{\tau}-\varepsilon_{\Delta}),\\ \nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta_{\rm b}\geq\\ \hskip 90.00014pt\ignorespaces-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-\epsilon_{\rm b}),\end{gathered}$
for all $\tau\in\{0,\Delta,\dots,T\}$.
@@
From Theorem 2, $\mathcal{C}_{\rm D}$ is controlled invariant, however the controlled invariance of $\mathcal{C}_{\rm I}$ itself cannot be proven without additional assumptions on $\boldsymbol{u}_{\rm b}$ and $\mathcal{C}_{\rm B}$. Therefore, feasibility of the optimization problem (DR-BCBF-QP) is not guaranteed. However, in the case that the optimization problem becomes infeasible, the robust backup control law $\boldsymbol{u}_{\rm b}$ can be used to stay in $\mathcal{C}_{\rm D}$ until the optimization problem becomes feasible again.
IV Numerical Results
To demonstrate the effectiveness of the proposed method at assuring safety under bounded disturbances, the results from two numerical examples are shown: a double integrator and a rigid body spacecraft with angular rate constraints.
@@ IV-A Double Integrator
$\dot{\boldsymbol{x}}=\begin{bmatrix}x_{2}\\ u\end{bmatrix}+\boldsymbol{d}_{x},$ (38)
@@
with a state vector $\boldsymbol{x}=[x_{1},x_{2}]^{T}\in\mathbb{R}^{2}$ where $x_{1}$ is the position and $x_{2}$ is the velocity, and an acceleration control variable $u\in\mathcal{U}=[-1,1]$. The safe set is defined as ${\mathcal{C}_{\rm S}\triangleq\{\boldsymbol{x}\in\mathbb{R}^{2}:-x_{1}\geq 0\}}$. $\boldsymbol{d}_{x}\in\mathcal{D}_{x}\subseteq\mathbb{R}^{n}$ is an unknown additive process disturbance and there exists a constant $\xi\in\mathbb{R}_{>0}$ such that $\left\|\boldsymbol{d}_{x}\right\|\leq\xi$. The backup control law $\boldsymbol{u}_{\rm b}(\boldsymbol{x})=-1$ will bring the system to the backup set ${\mathcal{C}_{\rm B}\triangleq\{\boldsymbol{x}\in\mathbb{R}^{2}:-x_{1}\geq 0,-x_{2}\geq 0\}}$ as long as $\xi<1$. In this particular example, the unknown disturbance is constant, with ${\boldsymbol{d}_{x}=\xi\frac{\boldsymbol{v}}{\left\|\boldsymbol{v}\right\|}}$ where ${\boldsymbol{v}=[1,1]^{T}}$, and ${\xi=0.08}$. The discretization time step used for computing the nominal backup flow is ${\Delta=0.02\mkern 4.0mu{\rm s}}$. The primary control law is ${\boldsymbol{u}_{\rm p}=1}$ which drives the system to the right half-plane (unsafe region).
@@
Using the proposed disturbance-robust backup CBF approach, the forward invariant set $\mathcal{C}_{\rm I}$ is computed for various values of $T$ between $0.5\mkern 4.0mu{\rm s}$ and $1.25\mkern 4.0mu{\rm s}$. These sets are plotted on Figure 2 in pink. The backup set, $\mathcal{C}_{\rm B}$, is plotted in blue and a robust controlled invariant set $\mathcal{C}_{\rm R}$ is plotted in green. For this simple linear system such a set can be computed by analytically solving for the flow and accounting for worst-case disturbances. The black dotted line represents the trajectory of the disturbed system using the proposed controller, with an integration horizon $T=1.25\mkern 4.0mu{\rm s}$. Notably, safety is maintained along this trajectory, as stated by Theorem 3. Nominal backup trajectories in gray emanate from the disturbed trajectory and the gray circles centered around them are the Gronwall norm balls from Lemma 2. The final Gronwall norm ball, colored in red, is always contained in $\mathcal{C}_{\rm B}$, as required by (20).
@@
The plots show that as the backup horizon $T$ is increased, the forward invariant set $\mathcal{C}_{\rm I}$ increases in size, up to a certain point. Since the Gronwall bound grows exponentially with time, the longer the backup horizon, the larger the final bound, $\delta_{\rm max}(T)$, will be. If $T$ is too large, the constraint on the terminal point of the nominal backup trajectory will dominate, and the set $\mathcal{C}_{\rm I}$ will begin to shrink. This therefore introduces a trade-off as $T$ cannot be made arbitrarily large when using Lemma 2. Naturally, as the disturbance bound $\xi$ increases, the size of $\mathcal{C}_{\rm I}$ will shrink since the bound on the backup flows is proportional to $\xi$.
@@
Figure 2: Phase space visualization of safety-critical control of double integrator system under bounded disturbances, using the proposed disturbance-robust backup control barrier function approach.

### v2

@@ III Disturbance Robustness
While the standard backup set method reviewed in Section II-B can guarantee safety for a system in which the dynamics are perfectly known, in practice there are always unmodeled parameters or external disturbances which perturb the dynamics. Therefore, it is desirable to leverage the advantages offered by the backup CBF approach, for dynamics with process disturbances. As such, consider the system
@@
We assume that a backup control law $\boldsymbol{u}_{\rm b}$ can be obtained which renders a backup set $\mathcal{C}_{\rm B}$ inside $\mathcal{C}_{\rm S}$ robustly forward invariant. This is made more precise below.
@@
Backup sets are often defined by a level set of a quadratic Lyapunov function based on the linearized dynamics about a stabilizable equilibrium point [3, 25], and a simple feedback controller such as a linear quadratic regulator can be used to render this set forward invariant. Techniques to robustify such quadratic Lyapunov functions have been studied in the literature [5],[26, Ch. 13.1],[27, Ch. 3]. While this robustification may result in a smaller backup set, this set is expanded to generate a larger controlled invariant set.
@@
[…]
@@
Let $\mathcal{L}_{h}$ and $\mathcal{L}_{h_{\rm b}}$ be the Lipschitz constants of $h$ and $h_{\rm b}$, respectively, and let $\delta_{{\rm max}}(\tau)$ be a norm bound on the deviation between $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ and $\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})$ at time $\tau\in[0,T]$:
$\left\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\right\|\leq\delta_{{\rm max}}(\tau),$ (19)
@@
These guarantee that $h_{\rm b}(\boldsymbol{\phi}^{d}_{\rm b}(T,\boldsymbol{x}))\geq 0$. Thus, all the functions which define $\mathcal{C}_{\rm D}$ are nonnegative, meaning that $\boldsymbol{x}\in\mathcal{C}_{\rm D}$ for all $\boldsymbol{x}\in\mathcal{C}_{\rm I}$, and so $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$. ∎
Lemma 1 assumes that a time-varying bound on the deviation between the nominal and disturbed backup flow, $\delta_{\rm max}(\tau)$, can be found. While problem-specific bounds can be obtained, we utilize a generalization of the Gronwall-Bellman inequality to obtain a bound for a wide class of nonlinear systems.
@@
For systems (4) and (12), let $f_{\rm cl}$ be locally Lipschitz on $\mathcal{X}$ with Lipschitz constant $\mathcal{L}_{\rm cl}$ and $\boldsymbol{d}_{x}$ be piecewise continuous in $\tau$ on $[0,T]$. If $\left\|\boldsymbol{d}_{x}\right\|\leq\xi$ for all $\boldsymbol{d}_{x}$ and some $\xi>0$, then for all $\tau\in[0,T]$ one has
$\left\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\right\|\leq\frac{\xi}{\mathcal{L}_{\rm cl}}(e^{\mathcal{L}_{\rm cl}\tau}-1)\triangleq\delta_{{\rm max}}(\tau).$
@@
Backup strategies often drive the system to an equilibrium and thus may be (at least weakly) contracting. When contraction bounds on the deviation between the disturbed and nominal backup flow can be obtained, $\delta_{{\rm max}}(\tau)$ can be made less conservative, and it will converge to a near-constant value as $T$ increases. Details on such bounds can be found in [28, Corollary 3.17]. A contraction bound is used effectively in the spacecraft rotation example in Section IV-B. For linear systems, flow deviation bounds can be even tighter.
Using the definition of $\mathcal{C}_{\rm D}$ and the corresponding robust backup controller $\boldsymbol{u}_{\rm b}$, we now examine the properties of $\mathcal{C}_{\rm D}$.
@@
Definition (15) with (22) and (23) completes the proof. ∎
@@
By Lemma 1, ${\boldsymbol{x}\in\mathcal{C}_{\rm I}\mkern-6.0mu\implies\boldsymbol{x}\in\mathcal{C}_{\rm D}}$, and by Lemma 3, $\boldsymbol{u}_{\rm b}$ ensures ${\boldsymbol{\phi}^{d}(\vartheta,\boldsymbol{x})\in\mathcal{C}_{\rm D}\subseteq\mathcal{C}_{\rm S}}$, ${\forall\mkern 2.0mu\vartheta\geq 0}$. ∎
We are now ready to establish the conditions that enable a controller to ensure the robust safety of (11). From the definition of $\mathcal{C}_{\rm I}$ we have
@@
Here, $\dot{\boldsymbol{x}}^{d}\triangleq f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}+\boldsymbol{d}_{x}$. Using this expansion, we can show that a controller which realizes forward invariance of $\mathcal{C}_{\rm I}$ keeps the disturbed system safe, and that conditions for such a controller can be directly computed, despite the unknown disturbance.
@@
As done in [5], the robustness terms $\eta$ and $\eta_{\rm b}$ upper-bound the unknown $\boldsymbol{d}_{x}$ term in (29). Thus the condition (34) implies (29). From a direct application of Theorem 1 to system (11), we obtain that (29) ensures ${\boldsymbol{\phi}^{d}(t,\boldsymbol{x}_{0})\in\mathcal{C}_{\rm I}}$, ${\forall t>0}$ for any ${\boldsymbol{x}_{0}\in\mathcal{C}_{\rm I}}$. From Lemma 1, we have $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$. ∎
Naturally, the original backup CBF constraints (8) are recovered in the absence of disturbances (i.e., $\xi=0$). As the constraints in (34) are continuous in $\tau$, the trajectory is again discretized similar to [4, Thm. 3] and appropriately tightened via a constant term $\varepsilon_{\Delta}$ where
@@
[…]
s.t. $\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta\geq$
$\hskip 80.00012pt\ignorespaces-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\epsilon_{\tau}-\varepsilon_{\Delta}),$
$\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta_{\rm b}\geq$
$\hskip 90.00014pt\ignorespaces-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-\epsilon_{\rm b}),$
for all $\tau\in\{0,\Delta,\dots,T\}$. As before, $\Delta\in\mathbb{R}_{>0}$ is a discretization time step satisfying $T/\Delta\in\mathbb{N}$. Because $\epsilon_{\tau}$ and $\epsilon_{\rm b}$ only depend on a priori known values and $\tau$, they can be pre-computed and reused each time (DR-bCBF-QP) is solved. Furthermore, the robustness terms $\eta$ and $\eta_{\rm b}$ depend on values that must be computed for the standard backup set method already, hence disturbance robustness adds negligible computational cost.
@@
From Theorem 2, $\mathcal{C}_{\rm D}$ is controlled invariant, however the controlled invariance of $\mathcal{C}_{\rm I}$ itself cannot be proven without additional assumptions on $\boldsymbol{u}_{\rm b}$ and $\mathcal{C}_{\rm B}$. Therefore, the feasibility of the optimization problem (DR-bCBF-QP) is not guaranteed. However, in the case that the optimization problem becomes infeasible, the robust backup control law $\boldsymbol{u}_{\rm b}$ can be used to stay in $\mathcal{C}_{\rm D}$ until the optimization problem becomes feasible again.
IV Numerical Examples
In this section we demonstrate the effectiveness of the proposed method in assuring safety under bounded disturbances using two simulation examples. Code and videos are available at: https://github.com/davidvwijk/DR-bCBF.
@@ IV-A Double Integrator
$\dot{\boldsymbol{x}}={\color[rgb]{0,0,0}\begin{bmatrix}x_{2},u\end{bmatrix}^{T}}\mkern-4.0mu+\boldsymbol{d}_{x},$ (36)
@@
with a state vector $\boldsymbol{x}=[x_{1},x_{2}]^{T}\in\mathbb{R}^{2}$ where $x_{1}$ is the position and $x_{2}$ is the velocity, and an acceleration control variable ${u\in\mathcal{U}=[-1,1]}$. The safe set is defined as ${\mathcal{C}_{\rm S}\triangleq\{\boldsymbol{x}\in\mathbb{R}^{2}:-x_{1}\geq 0\}}$. The unknown additive process disturbance is bounded with $\left\|\boldsymbol{d}_{x}\right\|\leq\xi\in\mathbb{R}_{>0}$. In this particular example, the unknown disturbance is constant, with ${\boldsymbol{d}_{x}=\xi\frac{\boldsymbol{v}}{\left\|\boldsymbol{v}\right\|}}$ where ${\boldsymbol{v}=[1,1]^{T}}$, and ${\xi=0.08}$. The backup control law ${\boldsymbol{u}_{\rm b}(\boldsymbol{x})=-1}$ brings the system to the backup set ${\mathcal{C}_{\rm B}\triangleq\{\boldsymbol{x}\in\mathbb{R}^{2}:-x_{1}\geq 0,-x_{2}\geq 0\}}$ as long as $\xi<1$. The discretization time step for computing the nominal backup flow is ${\Delta=0.02\mkern 4.0mu{\rm s}}$. The primary control law is ${\boldsymbol{u}_{\rm p}=1}$ which drives the system to the right half-plane (unsafe region).
@@
Using the proposed disturbance-robust backup CBF approach, the forward invariant set $\mathcal{C}_{\rm I}$ is computed for various values of $T$ between $0.5\mkern 4.0mu{\rm s}$ and $1.25\mkern 4.0mu{\rm s}$, plotted on Figure 2 in pink. The backup set, $\mathcal{C}_{\rm B}$, is plotted in blue and a robust controlled invariant set $\mathcal{C}_{\rm R}$ is plotted in green. For this simple linear system, such a set can be computed by analytically solving for the flow and accounting for worst-case disturbances. The black dotted line represents the trajectory of the disturbed system using the proposed controller, with an integration horizon $T=1.25\mkern 4.0mu{\rm s}$. Notably, safety is maintained along this trajectory, as stated by Theorem 3.
@@
The plots show that as the backup horizon $T$ is increased, the forward invariant set $\mathcal{C}_{\rm I}$ increases in size, up to a certain point. Since the Gronwall bound grows exponentially with time, the longer the backup horizon is, the larger the final bound, $\delta_{\rm max}(T)$, will be. If $T$ is too large, the constraint on the terminal point of the nominal backup trajectory will dominate, and the set $\mathcal{C}_{\rm I}$ will begin to shrink. This therefore introduces a trade-off as $T$ cannot be made arbitrarily large when using Lemma 2. Naturally, as the disturbance bound $\xi$ increases, the size of $\mathcal{C}_{\rm I}$ will shrink since the bound on the backup flows is proportional to $\xi$.
@@
Figure 2: Phase space visualization of safety-critical control for a double integrator system under bounded disturbances, using the proposed disturbance-robust backup control barrier function approach. Nominal backup trajectories in gray emanate from the disturbed trajectory (dotted black line) and the gray circles centered on the nominal trajectories are Gronwall norm balls from Lemma 2. The Gronwall norm ball at $\tau=T$, colored in red, is always contained in $\mathcal{C}_{\rm B}$, as required by (18).